OpenStudy (anonymous):

Help! f(x)=x^2-5 and g(x)=6x+2 (g x f)(x) The answer that I get is 6x^3+2x^2-5x+10 and I know it's wrong. What did I do wrong?

sam (.sam.):

(g x f)(x) gf(x) substitute f(x)=x^2-5 g(x^2-5) substitute g(x)=6x+2 6(x^2-5)+2

OpenStudy (anonymous):

ok where do i go from there?

OpenStudy (turingtest):

is it\[(g\circ f)(x)\]or\[(g\times f)(x)\]?

OpenStudy (anonymous):

the first one

OpenStudy (turingtest):

then .Sam. Has show you the right path; just simplify.

OpenStudy (anonymous):

right so 6x^3+2x^2-5x+10

OpenStudy (turingtest):

\[(g\circ f)(x)=g[f(x)]\]so what is g(x) ? write that out first

OpenStudy (anonymous):

(6x+2)(x^2-5)

OpenStudy (turingtest):

it is given in the problem that "f(x)=x^2-5 and g(x)=6x+2 " I just asked you what g(x) was...

OpenStudy (anonymous):

oh 6x+2

OpenStudy (turingtest):

cool, so now we want \(g[f(x)]\) which is just \(g(x)\) but replace all the x's that appear in g(x) with the expression for f(x)

OpenStudy (anonymous):

6(x^2-5)+2

OpenStudy (turingtest):

yes, now simplify :)

OpenStudy (anonymous):

6x^2-30+12

OpenStudy (turingtest):

where did you get 12 ?

OpenStudy (anonymous):

i mean 8

OpenStudy (turingtest):

...and where did you get the 8 from ?

OpenStudy (anonymous):

6+2

OpenStudy (turingtest):

where do you find 6+2 coming from 6(x^2-5)+2 ? watch your basic algebra, it'll bite you if you are not careful!

OpenStudy (anonymous):

6x^2-30+2

OpenStudy (anonymous):

i have no idea i hate math

OpenStudy (turingtest):

that's correct above; you only distribute the 6 to what is in the parentheses. of course we can still simplify this

OpenStudy (anonymous):

6x^2+32

OpenStudy (turingtest):

now you're just being lazy :P\[-30+2\neq32\]

OpenStudy (anonymous):

6x^2-28

OpenStudy (turingtest):

there we go :) I can feel your satisfaction from here

OpenStudy (anonymous):

bahaha yeah right. thank you!!

OpenStudy (turingtest):

welcome!